OpenStudy (calleyleeann):

TAYLOR EXPANSION: I have to find the first three terms of the Taylor expansion about x = 0 for f(x) = sin(x)/x. To get the first three terms I know I need to calculate up to the 4th derivative, which I did and verified by Mathematica. Here's what I know so far:

OpenStudy (loser66):

??

OpenStudy (calleyleeann):

\[f(x) = \frac{ \sin(x) }{ x }\] \[f'(x) = \frac{\cos\left(x\right)}{x}-\frac{\sin\left(x\right)}{x^2}\] \[f''(x)=\frac{2\sin\left(x\right)}{x^3}-\frac{\sin\left(x\right)}{x}-\frac{2\cos\left(x\right)}{x^2}\] \[f'''(x)=\frac{3\sin\left(x\right)}{x^2}+\frac{6\cos\left(x\right)}{x^3}-\frac{6\sin\left(x\right)}{x^4}-\frac{\cos\left(x\right)}{x}\] \[f''''(x)=-\frac{12\sin\left(x\right)}{x^3}+\frac{4\cos\left(x\right)}{x^2}-\frac{24\cos\left(x\right)}{x^4}+\frac{24\sin\left(x\right)}{x^5}+\frac{\sin\left(x\right)}{x}\]

OpenStudy (calleyleeann):

I was told to evaluate the limits, but for some of them I get a limit of infinity, which I can't really use...

OpenStudy (calleyleeann):

\[f(x_0+\epsilon)=f(x_0)+\epsilon f'(x_0)+\frac{ 1 }{ 2! }\epsilon^2 f''(x_0)+...\] This is what I'm using for the Taylor Expansion, where \[x_0=0\]

OpenStudy (calleyleeann):

DX

OpenStudy (calleyleeann):

Which one

OpenStudy (calleyleeann):

They seem very correct to me...

OpenStudy (calleyleeann):

But now what? I was told to do something with limits. The limit of f(x) is 1, which is the correct first term of the Taylor Expansion. However, when I do the limit for f'(x), I get infinity.... what do I do with that??

OpenStudy (calleyleeann):

I don't think I'm doing my limits right

OpenStudy (calleyleeann):

Calc 3

OpenStudy (calleyleeann):

I should probably know this, but how did you come up with the first part

OpenStudy (calleyleeann):

Nvm, makes sense

OpenStudy (calleyleeann):

Go on

OpenStudy (calleyleeann):

So regarding those first four terms of 1, 0, 0, 0, what do I do with those? Those aren't the first four terms of the taylor expansion

OpenStudy (calleyleeann):

So you're saying the taylor expansion is simply 1?

OpenStudy (loser66):

I think so. However, give me few minutes, I review my notes.

OpenStudy (calleyleeann):

Well, a quick search on Wolfram Alpha gives me an answer of \[1-\:\frac{x^2}{6}+\frac{x^4}{120}\] for the first three non-zero terms. Doesn't this mean the third term you found can't be zero?

OpenStudy (loser66):

oh, yes, it is, since \(lim_x\rightarrow 0 \dfrac{sinx}{x}=1\)

OpenStudy (loser66):

it is, if x =0

OpenStudy (loser66):

the all other terms but the first one =0 when x =0

OpenStudy (calleyleeann):

What?

OpenStudy (loser66):

What what? you have it, right? it just confirms our work.

OpenStudy (loser66):

see it, at the very first line of the theorem, it says lim (sinx/x) as x-->0 =1 http://planetmath.org/sites/default/files/texpdf/39255.pdf

OpenStudy (calleyleeann):

I know the lim of sin(x)/x is 1, but according to the taylor expansion calculated from Wolfram Alpha, the Taylor expansion is more than just 1 for the first couple orders. I understand what you did, but it doesn't come out to the "correct" answer. However, I don't understand how to get to the "correct" answer

OpenStudy (loser66):

you get \(sinx/x= 1-x^2/2!+x^4/5!-x^7/7!\), right? so, just plug x=0 in you have the first term is 1, second term is 0 and..... so on. the first three terms of it are 1, 0, 0 Dat sit

OpenStudy (calleyleeann):

Then I'm obviously missing something. How does that translate to the answer of \[1-\frac{ x^2 }{ 6 }+\frac{ x^4 }{ 120 }+...\]

OpenStudy (loser66):

hey!! 3! =6 and 5! =120

OpenStudy (calleyleeann):

Thank you I gotta go to class now

OpenStudy (loser66):

ok